Suppose x is a normally distrubuted random variable with a mean of 30 and standard deviation of 8. Find a value X0 of the random varible of such that. P(x>(Equal too) X)=0.5
1
Expert's answer
2021-10-27T14:04:06-0400
"P(X\geq X_0)=1-P(X<X_0)"

"=1-P(Z<\dfrac{X_0-\mu}{\sigma})=0.5"

"\dfrac{X_0-\mu}{\sigma}=0"

"X_0=\mu"

"X_0=30"